### v4

# SkillWrapper: Generative Predicate Invention for Task-level Planning

Abstract
Generalizing from individual skill executions to solving long-horizon tasks remains a core challenge in building autonomous robots. A promising direction is learning high-level, symbolic representations of the low-level skills of the robots, enabling reasoning and planning independent of the low-level state space. Recent advances in foundation models have made it possible to generate symbolic predicates that operate on raw sensory inputs—a process we call generative predicate invention—to facilitate downstream representation learning. However, it remains unclear which formal properties the learned representations must satisfy, and how they can be learned to guarantee these properties. In this paper, we address both questions by presenting a formal theory of generative predicate invention for task-level planning, resulting in symbolic operators useful for provably sound and complete planning. We propose SkillWrapper, a method that leverages foundation models to actively collect robot data and learn human-interpretable, plannable representations of black-box skills, using only RGB image observations. Our extensive empirical evaluation in simulation and on real robots shows that SkillWrapper learns abstract representations that enable solving unseen, long-horizon tasks in the real world with black-box skills.* Equal contribution. Work done primarily when the author was a postdoc researcher at Brown University.
I Introduction
An autonomous robot operating in the real world must process low-level sensory and motor signals while reasoning about high-level objectives [12, 33]. Analogous to how humans can perform complex tasks, like cooking or cleaning, without reasoning about muscle-level control, robots should have internal models of their skills that abstract away from the nuances of low-level control. Such models must capture the necessary conditions for executing a skill (e.g., “pouring from a teapot requires holding it first”) and the consequences of doing so (e.g., “pouring from a teapot makes it empty”). These two properties, known as preconditions and effects in the AI planning literature [18], enable compositional reasoning to identify long-horizon plans that can sequence lower-level skills to solve a task. Typically, these models must be specified manually. However, in real-world settings, such task-level models may be nontrivial to acquire due to complex inter-skill constraints specific to the robot’s embodiment. This calls for algorithms that autonomously learn symbolic transition models of black-box skills, enabling robots to solve long-horizon tasks using off-the-shelf AI planners.
Traditional approaches for learning task-level models often require either factorizing the low-level state space to learn classifiers for each symbolic representation [35], privileged information such as object poses [53], or initial pool of task-level symbols [57] (detailed in Appendix -F). Foundation models have enabled a new paradigm: generating semantically meaningful predicates from raw observations and directly evaluating their truth values on low-level observations (e.g., RGB images)—a process we refer to as generative predicate invention. While recent work has explored how foundation models can be used for predicate invention by generating Python code to implement predicates [39] or undirected sampling followed by a hill-climbing model search [4], these methods do not provide any concrete guarantees of correctness, leaving core questions unanswered: what properties should these learned abstractions satisfy, and how can they be learned to achieve these properties?
Our answer is twofold. First, we develop a formal theory of generative predicate invention for task-level planning, precisely characterizing the conditions under which a learned model is provably sound and complete with respect to downstream planning. Building on this foundation, we introduce SkillWrapper, a method explicitly designed to guarantee these theoretical criteria. SkillWrapper uses foundation models in three ways: interactively collecting data in the robot’s environment, proposing predicates when the current model fails, and classifying predicate truth values based solely on RGB image observations. Using these data and predicates, SkillWrapper learns symbolic representations of black-box skills that are both human-interpretable and directly usable for task-level planning. To the best of our knowledge, SkillWrapper is the first work integrating active data collection and generative predicate invention using foundation models while provably guaranteeing soundness and completeness for task-level planning.
We highlight the following contributions: (1) A formal theory of generative predicate invention for provably sound and complete representations; (2) SkillWrapper, a principled system built on this framework that leverages foundation models to learn interpretable symbolic representations of black-box skills; and (3) an extensive empirical evaluation of the system in simulated and real robotic platforms.
Figure 1: Overview of  SkillWrapper. For a robot equipped with black-box skills, SkillWrapper learns task-level plans that are compatible with off-the-shelf planners. These representations are comprised of predicates invented by the foundation model. Given a novel planning problem described using the initial state and goal state as RGB images, a foundation model produces the corresponding abstract states by applying the invented predicates to the low-level states. SkillWrapper is agnostic to the robot embodiment, and we illustrate both real-world robots and simulation in this figure.
II Formal Framework
We consider a problem setting in which a robot is equipped with a set of predefined, “black-box” skills. The robot can evaluate whether a skill is executable in the current state, but it does not have a complete transition model of the skills a priori, and therefore cannot compose its skills to solve long-horizon problems without considering low-level details. However, if the robot were to learn a symbolic model of its skills, it could use classical task planning approaches to efficiently compose its skills and enable the robot to solve novel tasks. We formalize this setting as the problem of Skill Model Learning.
II-A Preliminaries
Environment model
We define an environment as a tuple $({\mathcal{S},\mathcal{T},\Omega,T})$, where the state space $\mathcal{S}$ is assumed to be high-dimensional, continuous, and fully observable; the type set $\mathcal{T}$ enumerates the possible object types; and the skills $\Omega$ are object-centric, such that each $\omega\in\Omega$ is parameterized by object types drawn from the type set $\mathcal{T}$. The transition function $T:\mathcal{S}\times\Omega\to\mathcal{S}$ characterizes the environment dynamics but is unknown to the robot. For example, Pour(?teapot, ?mug) may be used to pour tea from a teapot into a mug. However, certain environmental aspects (e.g., which mugs are available to pour into) may differ between settings during learning. We therefore define a specific setting by the tuple $({{s}_{0},\mathcal{O}})$, specifying an initial state ${s}_{0}\in\mathcal{S}$ and a set of typed objects $\mathcal{O}$, where the type(s) of object $o\in\mathcal{O}$ are denoted by $\texttt{t}_{o}\subseteq\mathcal{T}$.
Black-box skills
We model skills as object-centric options [60] with discrete, object-typed parameters. Formally, a skill $\omega\in\Omega$ is defined by a tuple $({\mathcal{I}_{\omega},\pi_{\omega},\beta_{\omega},\Theta_{\omega}})$, where the initiation set $\mathcal{I}_{\omega}\subseteq\mathcal{S}$ contains states from which the skill may be executed; the policy $\pi_{\omega}$ controls the robot during the skill; the termination set $\beta_{\omega}\subseteq\mathcal{S}$ is the set of states at which the skill immediately terminates; and the skill parameters $\Theta_{\omega}=({\theta_{\omega}^{1},\dots,\theta_{\omega}^{k}})$ specify type constraints $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\mathcal{T}$ for valid skill arguments. Specifically, a skill $\omega\in\Omega$ may be instantiated to create a skill instance $\underline{\omega}$ using objects ${o}_{1:{k}}\in\mathcal{O}^{k}$ if and only if for $1\leq i\leq k$, $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\texttt{t}_{o_{i}}$. We assume that executing a skill only affects the state of objects passed as arguments. [^1] We do not, however, assume that a skill necessarily affects the state of all objects passed in as arguments.
Symbolic abstractions
Because the robot must evaluate $\mathcal{I}_{\omega}$ on individual states, it cannot distinguish essential skill information from irrelevant details, making long-horizon planning combinatorially difficult. However, each skill affects only a few objects at once, leaving the rest of the world unchanged. The robot can exploit this property by instead learning a factored state representation—formalized here as first-order logic models in PDDL [43]—providing an abstract transition model of its skills. Such abstractions enable the robot to use classical task planning to compose its skills and accomplish unseen, long-horizon goals.
We use symbolic predicates $\mathcal{P}$ to express abstract relations between objects. Each predicate $\sigma\in\mathcal{P}$ is a tuple $({C_{\sigma},\Theta_{\sigma}})$, where the predicate classifier $C_{\sigma}$ tests whether the predicate holds in a state, given a binding of objects to the predicate parameters $\Theta_{\sigma}=({\theta_{\sigma}^{1},\dots,\theta_{\sigma}^{n}})$. Each parameter $\theta_{\sigma}^{i}$ specifies a type constraint $\texttt{t}_{\theta_{\sigma}^{i}}\subseteq\mathcal{T}$ on corresponding object arguments. Formally, the partial function $C_{\sigma}:\mathcal{O}^{n}\rightharpoonup(\mathcal{S}\to\{0,1\})$ is defined on objects ${o}_{1:{n}}\in\mathcal{O}^{n}$ if and only if $\texttt{t}_{\theta_{\sigma}^{i}}\subseteq\texttt{t}_{o_{i}}$ for $1\leq i\leq n$. In such cases, we say that the objects ${o}_{1:{n}}$ are valid arguments for $\sigma$.
By observing different effects, the robot must construct a set of operators $\mathcal{A}$ that define an abstract transition model for the robot’s skills. We define each operator $a\in\mathcal{A}$ by $({\omega_{a},\Theta_{a},\textsc{Pre},\textsc{Eff}^{+},\textsc{Eff}^{-}})$, where $\omega_{a}\in\Omega$ is the corresponding skill; the operator parameters $\Theta_{a}=({\theta_{a}^{1},\dots,\theta_{a}^{m}})$ impose type constraints $\texttt{t}_{\theta_{a}^{i}}\subseteq\mathcal{T}$ on operator arguments; the preconditions Pre are a conjunction of literals over $\mathcal{P}$ defining the conditions necessary to apply the operator; and the add and delete effects, $\textsc{Eff}^{+}$ and $\textsc{Eff}^{-}$, are the subsets of $\mathcal{P}$ that become true and false, respectively, after the operator is applied.
Grounded Abstractions
To apply an abstract transition model to a real-world setting, the robot must map the low-level state to a high-level abstract state, which requires grounding the known predicates using concrete objects so that their truth value may be determined. A predicate $\sigma\in\mathcal{P}$ may only be grounded using valid arguments, inducing a grounded predicate $\underline{\sigma}=\sigma({o}_{1:{n}})$. Given a low-level state, the classifier $C_{\underline{\sigma}}:\mathcal{S}\to\{0,1\}$ tests whether the predicate holds for the object arguments. We define the set of grounded predicates for a setting $({{s}_{0},\mathcal{O}})$ under predicates $\mathcal{P}$ as $\underline{\mathcal{P}}=\{\sigma({o}_{1:{n}})\,:\,\sigma\in\mathcal{P},{o}_{1:{n}}\in\mathcal{O}^{n},\bigwedge_{i=1}^{n}\texttt{t}_{\theta_{\sigma}^{i}}\subseteq\texttt{t}_{o_{i}}\}$, inducing an abstract state space $\underline{\mathcal{S}}=2^{\underline{\mathcal{P}}}$ where each abstract state corresponds to a specific set of grounded relations.
Given a set of operators $\mathcal{A}$ in a setting $({{s}_{0},\mathcal{O}})$, the abstract action space $\underline{\mathcal{A}}$ is the set of all valid groundings of the operators using objects from $\mathcal{O}$. Each abstract action (i.e., grounded operator) is defined as $\underline{a}=a({o}_{1:{m}})$, where $a\in\mathcal{A}$, ${o}_{1:{m}}\in\mathcal{O}^{m}$, and for $1\leq i\leq m$, $\texttt{t}_{\theta_{a}^{i}}\subseteq\texttt{t}_{o_{i}}$. Grounding an operator induces ground preconditions $\underline{\textsc{Pre}}$, which are a conjunction of ground literals over $\underline{\mathcal{P}}$; ground add effects $\underline{\textsc{Eff}}^{+}\subseteq\underline{\mathcal{P}}$; and ground delete effects $\underline{\textsc{Eff}}^{-}\subseteq\underline{\mathcal{P}}$.
The grounding function $\mathcal{G}:\underline{\mathcal{P}}\to 2^{\mathcal{S}}$ maps each grounded predicate to its grounding set $\mathcal{G}({\underline{\sigma}})\subseteq\mathcal{S}$, defined as $\mathcal{G}({\underline{\sigma}})=\{{s\in\mathcal{S}\,:\,C_{\underline{\sigma}}(s)=1}\}$. We overload this notation for abstract states $\underline{s}\in\underline{\mathcal{S}}$ so that $\mathcal{G}({\underline{s}})=\bigcap_{\underline{\sigma}_{i}\in\underline{s}}\mathcal{G}({\underline{\sigma}_{i}})$. Conversely, the abstraction function $\textsc{Abstract}:\mathcal{S}\to\underline{\mathcal{S}}$ maps each low-level state $s\in\mathcal{S}$ to the abstract state $\underline{s}\in\underline{\mathcal{S}}$ defined by $\textsc{Abstract}(s)=\{{\underline{\sigma}\in\underline{\mathcal{P}}\,:\,C_{\underline{\sigma}}(s)=1}\}$.
II-B Problem Definition
We now define the task-level model learning problem and a task-level solution as follows:
Definition 1.
Given an environment $({\mathcal{S},\mathcal{T},\Omega,T})$ containing settings $\{({{s}_{0},\mathcal{O}})\}_{i=1}^{N}$, we define a Task-Level Model Learning problem as learning an abstract transition model $\mathcal{M}=({\mathcal{P},\mathcal{A}})$ for the skills $\Omega$.
After a period of continual learning in one or more settings, a robot may be evaluated on a task-level planning problem $\mathbf{p}=({{s}_{0},\mathcal{O},\mathcal{S}_{g}})$, where $\mathcal{S}_{g}\subseteq\mathcal{S}$ is the set of goal states to be reached. Given a model $\mathcal{M}=({\mathcal{P},\mathcal{A}})$, a classical planner can be used to search for an abstract plan $\underline{a}_{1},\dots,\underline{a}_{n}$}] that solves the abstract planning problem $\underline{\mathbf{p}}=({{\underline{s}}_{0},\mathcal{O},\underline{\mathcal{S}}_{g}})$.
Definition 2.
An abstract plan $\underline{a}_{1},\dots,\underline{a}_{n}$}] is a solution for skill planning problem $\mathbf{p}=({{s}_{0},\mathcal{O},\mathcal{S}_{g}})$ iff for $1\leq i\leq n,\,s_{i}=T(s_{i-1},\underline{\omega}_{i}),s_{i-1}\in\mathcal{I}_{\underline{\omega}_{i}},\text{ and }s_{n}\in\mathcal{S}_{g},$ where $\underline{\omega}_{i}=\underline{\omega}_{\underline{a}_{i}}$.
In Sec. III, we describe how SkillWrapper constructs $\mathcal{M}$} from raw skill executions.
III SkillWrapper
In this section, we introduce SkillWrapper, a novel approach that autonomously learns symbolic representations for black-box skills while providing several guarantees such as soundness and completeness (see Appendix -B for further theoretical details). To produce a valid abstract model that enables planning,  SkillWrapper iterates through a three-step process: 1) actively proposing and executing exploratory skill sequences to collect data on the initiation and termination set of each skill, 2) incrementally building a set of predicates from scratch by contrasting positive and negative examples, and then 3) constructing valid operators using these invented predicates, from which further exploratory skill sequences can be proposed. This procedure is outlined in Algorithm 1. As SkillWrapper continues to collect data, invent and incorporate new predicates, and update its planning model, it progressively improves the accuracy of the learned task-level transition model. The resulting task-level representations, or operators, can be used with an off-the-shelf classical planner to solve task planning problems specified using RGB images. We delve into the core components of our system (lines 4–9 of Algorithm 1) in the following subsections. Importantly, we provide strong theoretical results for the soundness and completeness of SkillWrapper in Section III-D.
Algorithm 1 SkillWrapper
1:Input: Set of skills $\Omega$, number of iterations $m\in\mathbb{N}_{1}$
2:Output: Abstract transition model $\mathcal{M}=(\mathcal{P},\mathcal{A})$
[…]
4:for $i\in\{1,\dots,m\}$ do
5:  $\langle\underline{\omega}_{1},\ldots,\underline{\omega}_{k}\rangle\leftarrow\textsc{ProposeSkillSequence}(\Omega,\mathcal{D},\mathcal{P},\mathcal{A})$
6:  $\mathcal{D}\leftarrow\mathcal{D}\,\|\,\textsc{ExecuteSkills}(\underline{\omega}_{1},\ldots,\underline{\omega}_{k})$
7:  $\mathcal{P}\leftarrow\textsc{InventPredicates}(\Omega,\mathcal{D},\mathcal{P},\mathcal{A})$
8:  $\mathcal{A}\leftarrow\textsc{LearnOperators}(\mathcal{D},\mathcal{P})$
9:return $\mathcal{M}$
III-A Active Data Collection
To collect data for learning, our method computes a sequence of skills $\langle\underline{\omega}_{1},\dots,\underline{\omega}_{k}\rangle$ for the robot to execute and collects the resulting transitions. While executing these skill sequences, the robot collects a dataset $\mathcal{D}$ of transitions of the form $\langle s,\underline{\omega},s^{\prime}\rangle$, where $s,s^{\prime}\in{\mathcal{S}}$. These transitions can answer two questions:
  1. 1.
    Executability: Can the skill instance $\underline{\omega}$ be executed from state $s$?
  2. 2.
    Skill Dynamics: If $s\neq s^{\prime}$, what has changed in the environment due to executing $\underline{\omega}$?
We guide the exploration of skill preconditions and effects using a foundation model, which proposes skill sequences in natural language. Rather than naively sampling these sequences from the foundation model’s token distribution, we prompt the model for a batch of candidate skill sequences and apply two scoring functions to bias the system toward promising sequences that explore novel skill instance pairs and keep a balance between success and failure executions. SkillWrapper facilitates the efficient data collection process, which results in a dataset $\mathcal{D}$ that is critical to downstream processes, such as guiding predicate invention with failure transitions, and improving the learned abstract model by eliminating unnecessary preconditions, etc.
We prompt a foundation model with the robot’s skill set $\Omega$ and the current set of abstract predicates $\mathcal{P}$ to generate and propose sequences of skills $\sigma$, with which the robot collects a dataset of transitions $\mathcal{D}$. SkillWrapper assigns a score (Alg. 2) tuple $(C,Ch)$ to all sequences and maintains a subset of pareto-front sequences that cannot strictly dominate another sequence, i.e., $(C_{i}<C_{j})\lor(Ch_{i}<Ch_{j})\lor(C_{i}\leq C_{j}\land Ch_{i}\leq Ch_{j})$ where $i\neq j$. An output skill sequence is finally chosen from this Pareto-front subset (see App. -A1 for more details).
Coverage ($C$)
Coverage evaluates the information gain on all possible pairs of consecutively executed skills over existing transitions after executing a new skill sequence. Specifically, the information gain is measured by the increase in Shannon entropy [54] over the distribution of all consecutive skill pairs resulting from executing the proposed skill sequence
$C=\frac{\mathcal{Q}^{{}^{\prime}}}{\Sigma\mathcal{Q}^{{}^{\prime}}}\times log(\frac{\mathcal{Q}^{{}^{\prime}}}{\Sigma\mathcal{Q}^{{}^{\prime}}})-\frac{\mathcal{Q}}{\Sigma\mathcal{Q}}\times log(\frac{\mathcal{Q}}{\Sigma\mathcal{Q}})$
where $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are matrices that tabulate the number of pairs of consecutively executed skills that occur before and after executing a new skill sequence. Maximizing coverage encourages the generation of proposed skill sequences that contain the least explored skill pairs. More importantly, this allows SkillWrapper to uncover a larger set of interdependencies across the preconditions and effects of all skills.
Chainability ($Ch$)
Chainability predicts the ratio of successful to failed pairs of consecutively executed skills. By computing chainability, we estimate the degree to which the preconditions of operators learned in each iteration are satisfied, and executability can be inferred from the estimated symbolic states and the operators. With an appropriate chainability score, the collected dataset of skill execution traces maintains a balance between the number of successful executions and failure executions, which is ideal for identifying possible mismatched pairs and thus inventing predicates.
Algorithm 2 Propose Skill Sequences
1:Input: Skill set $\Omega$, skill execution traces $\mathcal{D}$, predicate set $\mathcal{P}$, operator set $\mathcal{A}$, batch size $n$
2:Output: Proposed skill sequence $\sigma$
[…]
5:for $\langle\underline{\omega}_{1},\ldots,\underline{\omega}_{k}\rangle$ in $\mathrm{seq\_batch}$ do
6:  $\mathrm{cov}\leftarrow\textsc{Coverage}(\mathcal{D},\langle\underline{\omega}_{1},\ldots,\underline{\omega}_{k}\rangle)$
7:  $\mathrm{chain}\leftarrow\textsc{Chainability}(\mathcal{A},\langle\underline{\omega}_{1},\ldots,\underline{\omega}_{k}\rangle)$
8:  $\mathrm{Scores}[\langle\underline{\omega}_{1},\ldots,\underline{\omega}_{k}\rangle]\leftarrow(\mathrm{cov},\mathrm{chain})$
9:return $\textsc{ParetoOptimal}(\mathrm{Scores})$
III-B Predicate Invention
We now present our predicate invention algorithm. Unlike previous work [57, 39], our algorithm does not require an initial set of predicates or partially complete operators to bootstrap the invention process.
Conditions for predicate invention
SkillWrapper identifies two conditions under which the current predicate vocabulary is insufficient, based on the desired properties of soundness and completeness (discussed fully in Appendix -B). In these cases, the system must invent new predicates to resolve incongruities between the observed data and the current abstract transition model. Intuitively, SkillWrapper only introduces new predicates if and only if they are necessary to preserve the soundness and completeness of the model given the new transition. This is explained in detail in Appendix -D.
Algorithm 3 Invent Predicates
1:Input: Skill set $\Omega$, skill execution traces $\mathcal{D}(\omega)=\{\langle\mathrm{s},\underline{\omega},\mathrm{s}^{\prime}\rangle\}_{\omega}$, predicate set $\mathcal{P}$, operator set $\mathcal{A}$.
2:Output: Predicate set $\mathcal{P}$
3:for $\omega\in\Omega$ do
4:  while precondition inconsistency exists do $\triangleright$ equation 1
5:    $\sigma\leftarrow\textsc{NewPredicate}$
6:    $\mathcal{P}\leftarrow\mathcal{P}\,\|\,\sigma$ if ScorePrecond($\sigma,\mathcal{P},\omega,\mathcal{D}$)
7:  while effect inconsistency exists do $\triangleright$ equation 2
8:    $\sigma\leftarrow\textsc{NewPredicate}$
9:    $\mathcal{P}\leftarrow\mathcal{P}\,\|\,\sigma$ if ScoreEff($\sigma,\mathcal{P},\omega,\mathcal{D}$)
10:return $\mathcal{P}$
To formally describe the conditions, we define two sets, $\bm{\alpha}_{\omega}$ and $\bm{\zeta}_{\omega}$, representing the states in which the model predicts that the skill may either be initiated (when $s\in\bm{\alpha}_{\omega}$) or terminated (when $s\in\bm{\zeta}_{\omega}$), respectively. Both sets are derived from and defined with the learned operators, and their formal definitions can be found in Appendix -B.
The first condition arises when SkillWrapper detects that the symbolic vocabulary cannot express a necessary precondition for a skill. Concretely, this occurs when two transitions involve instances of the same skill, one successful and one failed, both satisfy the initiation condition of the skill under the current predicates. Because the vocabulary cannot distinguish between these initial states, an additional predicate is required. Formally, this condition is expressed as:
$\begin{split}\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in\mathcal{D}\text{ s.t. }s_{i}\in\bm{\alpha}_{\underline{\omega}_{i}},s_{j}\in\bm{\alpha}_{\underline{\omega}_{j}},\\ \text{but }s_{i}\in\mathcal{I}_{\underline{\omega}_{i}}\text{ and }s_{j}\notin\mathcal{I}_{\underline{\omega}_{j}}.\end{split}$ (1)
The second condition used by SkillWrapper to trigger predicate invention is based on inconsistencies in observed skill effects. Specifically, this occurs when two transitions that involve instances of the same skill produce identical abstract effects, despite one succeeding and the other failing. In a deterministic setting, this reduces to a successful skill execution producing no effects, though the condition naturally extends to stochastic settings with mid-execution failures. Formally, we express the condition as:
Figure 2: Example of Predicate Invention. The initial states of two transitions are both said to satisfy the preconditions of certain operators learned from the same skill, while transition 1 is successful, but transition 2 is not. In this case, the first condition (equation 1) is triggered, and the foundation model is prompted with both transitions to invent a new predicate.
$\begin{split}\exists\langle s_{i},\underline{\omega}_{i},s^{\prime}_{i}\rangle,\langle s_{j},\underline{\omega}_{j},s^{\prime}_{j}\rangle\in{\mathcal{D}}\text{ s.t. }s_{i}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{i}},s_{j}^{\prime}\in\bm{\zeta}_{\underline{\omega}_{j}},\\ \text{ but }s_{i}\in{\mathcal{I}}_{\underline{\omega}_{i}}\text{ and }s_{j}\notin{\mathcal{I}}_{\underline{\omega}_{j}}.\end{split}$ (2)
Contrastive predicate proposal
When a satisfying transition pair is identified under the conditions of predicate invention, SkillWrapper prompts the foundation model with the two transitions and their corresponding states (RGB images) to propose a candidate predicate that can potentially distinguish the transition pair. The transition pair offers contrastive clues for the foundation model to propose predicates that precisely resolve the incongruity.
Empirical predicate selection
Although foundation models provide a strong prior on which information is skill-relevant for predicate proposal, they may produce errors and hallucinations. To ensure robustness, we introduce a scoring function that estimates the usefulness of a candidate predicate by trying adding it to the vocabulary and learning hypothetical operators. For each successful transition in $\mathcal{D}$}, there must exist at least one operator with preconditions satisfied by the transition’s initial state; for each failed transition, no such operator may exist. Effect evaluation follows the same principle. After evaluating over $\mathcal{D}$, we decide whether to add the candidate predicate based on an empirical threshold. Further details of the scoring function can be found in Algorithm 6 (App. -A).
III-C Operator Learning
Our operator learning procedure extends the associative model learning paradigm [3] to the setting of skill abstraction.
Associative model learning
A single skill may induce multiple distinct abstract state changes depending on the context of execution. To represent these conditional effects, SkillWrapper clusters observed transitions based on their lifted (object-agnostic) effect sets, enabling it to learn a single operator across distinct instantiations of a skill. The preconditions of each operator are then computed as the intersection of all initial abstract states in the corresponding transitions, ensuring that each operator is both minimal and consistent. The learning process operates directly on the abstract level, allowing it to capture relationships over objects beyond the skill parameters. We provide a detailed explanation in App. -G.
Multi-type object-centrism
In realistic domains, objects do not fit neatly into single type categories, but rather belong to multiple overlapping categories (e.g., a $\mathtt{Cup}$ is $\mathtt{fillable}$, while a $\mathtt{Bottle}$ is both $\mathtt{fillable}$ and $\mathtt{openable}$). This complicates the process of generalizing grounded transitions into lifted operators, because it may be ambiguous which object attribute enables successful execution. We adopt a conservative strategy: SkillWrapper assigns arguments of each operator using the lowest level of the type hierarchy consistent with the data, preventing over-generalization while retaining compositional structure.
Predicate re-scoring
Predicates are generated sequentially, and early inventions may bias later stages if left unchecked. To mitigate this, SkillWrapper re-applies the scoring function to the entire predicate set after each iteration of data collection. This allows spurious or redundant predicates to be discarded as more data is collected. In addition, tautological predicates—those that are always true or always false—are automatically filtered. As a result, the learned predicate set remains compact, informative, and aligned with the most recent transition data.
III-D Theoretical Analysis
We now provide theoretical guarantees for SkillWrapper. These results show that the learned symbolic model is sound with respect to observed data and converges to a complete model with high probability. Full proofs are deferred to App. -C.
Theorem 1 (Soundness of SkillWrapper).
Every operator $a\in{\mathcal{A}}_{n}$ in the model ${\mathcal{M}}_{n}$ learned by SkillWrapper is supported by at least one observed transition $\langle s,\omega,s^{\prime}\rangle\in B_{n}$. That is, $\underline{s}\models\textsc{Pre}_{a}$ and $\underline{s}^{\prime}\models\textsc{Eff}_{a}$.
Sketch. SkillWrapper constructs each operator directly from sampled transitions; thus no unsupported operator can appear.
Lemma 1.
For each $\omega\in\Omega$, the initiation set $I_{\omega}$ and termination set $\bm{\beta}_{\omega}$ inferred by SkillWrapper match exactly the corresponding predicate sets $\bm{\alpha}_{\omega}$ and $\bm{\zeta}_{\omega}$ derived from $B_{n}$, i.e. $I_{\omega}=\bm{\alpha}_{\omega}$ and $\bm{\beta}_{\omega}=\bm{\zeta}_{\omega}$.
Sketch. Any mismatch would contradict the termination condition of SkillWrapper; thus initiation and termination sets are consistent with observed data.
Theorem 2 (Probabilistic-completeness of SkillWrapper).
Let $\mathcal{M}$$^{*}$ be the true complete model. With probability at least $1-|{\mathcal{H}}|\exp(-n\epsilon)$, the model $\widehat{{\mathcal{M}}}_{n}$ learned from $n$ i.i.d. samples satisfies $d_{\mathrm{compl}}(\widehat{{\mathcal{M}}}_{n},{\mathcal{M}}^{*})\leq\epsilon$, i.e. it misses fewer than an $\epsilon$-fraction of feasible transitions.
Sketch. Since $\widehat{\mathrm{Err}}(\widehat{{\mathcal{M}}}_{n})=0$ by construction, a Chernoff bound and union bound over the finite hypothesis class imply that the true error is small with high probability. (We derive the upper bound of the hypothesis class in Appendix -E.)
Together, these results establish that SkillWrapper learns symbolic operators that are sound with respect to observed transitions, consistent in their preconditions and effects, and probabilistically complete relative to the true underlying model. These properties justify SkillWrapper as a reliable model-learning procedure for planning. Next, we discuss empirical evaluation for SkillWrapper.
IV Experiments
For all experiments in this section, we consider images as fully observable state representations, assuming that an abstract state can be inferred from an image without uncertainty. Both the initial and goal states of each problem are specified using RGB images. These images may come from diverse sources, including a top-down view of an animated game, a third-person camera observing a robot, or the robot’s own egocentric perspective. All quantitative results reported in this section are averaged over five independent runs for simulation experiments and three runs for real robot experiments.
(a) Franka: Initial
(b) Franka: Goal
(c) Bimanual Kuka: Initial
(d) Bimanual Kuka: Goal
Figure 3: Initial and goal states for robot experiments.
IV-A Implementation of Predicates
We employ foundation models (specifically vision-language models or VLMs) for both predicate generation and evaluation:
  • A foundation model gives us a string that can be used as a lifted predicate (in other words, it generates interpretable relational classifiers with a good heuristic).
  • After grounding with valid parameters, the predicate can be prompted to the foundation model again to acquire the truth value. In other words, a foundation model can be used as a relational classifier.
With these two properties, we can use the VLM’s response in string form as a relational predicate, and a grounded version of the predicate can be used as a classifier. We use GPT-5 [47] for predicate generation and evaluation. In addition to the system performance reported in the main paper, we also conducted comprehensive studies of the component-wise reliability of the VLM in Appendix -I, including predicate invention and classification tests, per-predicate accuracy, and robustness to lighting conditions and view angle changes. This proves the real-world robustness of SkillWrapper.
(a) Robotouille: Initial
(b) Robotouille: Goal
Figure 4: Initial and goal state images in the Robotouille domain.
IV-B Simulation
We first conduct experiments in Robotouille [19], a simulated grid-world kitchen domain with a robot that has five high-level skills: Pick, Place, Cut, Cook, and Stack. In the environment, there are several objects: a patty, lettuce, a top bun, and a bottom bun; there is also a cutting board and a stove for cutting the lettuce and cooking the patty, respectively. We design and categorize 50 abstract planning problems: 20 easy problems, whose optimal solutions have no more than 7 steps; 20 hard problems, whose optimal solutions have no more than 15 steps; and 10 impossible tasks that cannot be realized in the environment, which a sound model must identify.
For a complete evaluation of SkillWrapper, we compare SkillWrapper against the following baselines:
  • Expert Operators: A human expert familiar with PDDL is asked to interact with the environment and manually write predicates with semantics and PDDL operators.
  • System Predicates: This baseline directly uses the built-in predicate set of the simulator, which is designed to define any possible simulated state unambiguously. In addition, instead of evaluating the predicate truth values with foundation models, this baseline has access to the exact abstract states of the simulator. However, predicate invention is disabled in this environment.
  • ViLa [24]: This baseline is a closed-loop VLM-based approach that iteratively prompts a foundation model for the next action until the goal state is reached, given an image observation and the robot’s action history.
  • Random Exploration: Instead of proposing skill sequences, this baseline, acting as a lower-bound, randomly samples a skill and populates the arguments with valid objects. This baseline shares the same predicate invention and operator learning algorithms as SkillWrapper.
  • FM Invent: This baseline invents predicates in batches by directly prompting the foundation model with all image states, instead of using the predicate invention method of SkillWrapper. The active data collection and operator learning procedures are the same as SkillWrapper.
  • No Heuristic: This baseline is the same as SkillWrapper, except that skill sequences are selected randomly from the foundation model’s output without applying the heuristics (Sec. III-A).
For each baseline that performs operator learning, we run the learning algorithm for five iterations, with each iteration proposing and executing one skill sequence comprised of 15 steps as its interaction budget. We then evaluate each method on the evaluation set and report the average results in Table I, where Solved % is the percentage of the problem set that was successfully solved or where impossible tasks were correctly identified by returning an empty plan, and PB stands for planning budgets—the number of plans that were tried before solving the problem (adopted from [39]). We set a planning budget cap of 10 across all problems; if the planning budget has been used up for a problem, we consider it a failure. Theoretically, PB is an adequate metric that reflects the completeness of the learned model, and the impossible problems reflect its soundness. As shown in the table, SkillWrapper outperforms all baselines that have no access to privileged knowledge, and even surpasses the performance of the System Predicates baseline on hard problems. For brevity, we leave case studies, example operators, failure modes, and analysis of simulated results in Appendix -H.
TABLE I: Baseline Comparison in Robotouille Environment
Method Easy Hard Impossible Solved % $\uparrow$ PB $\downarrow$ Solved % $\uparrow$ PB $\downarrow$ Expert Ops. 81.0 $\pm$ 3.7 1.9 $\pm$ 0.4 58.1 $\pm$ 3.9 4.2 $\pm$ 0.4 100 $\pm$ 0.0 Sys Preds. 79.0 $\pm$ 3.7 2.6 $\pm$ 0.2 22.0 $\pm$ 12.9 7.8 $\pm$ 1.3 42.0 $\pm$ 7.5 ViLa [24] 46.0 $\pm$ 16.2 - 13.9 $\pm$ 11.6 - 20.0 $\pm$ 10.9 Random Exp. 4.0 $\pm$ 2.0 9.6 $\pm$ 0.2 0 $\pm$ 0.0 10.0 $\pm$ 0.0 100 $\pm$ 0.0 FM Invent 23.3 $\pm$ 16.5 7.7 $\pm$ 1.7 1.7 $\pm$ 2.4 9.7 $\pm$ 0.2 63.3 $\pm$ 26.2 No Heuristic 76.0 $\pm$ 4.9 2.5 $\pm$ 0.9 24.0 $\pm$ 19.6 7.8 $\pm$ 1.8 80 $\pm$ 20.9 Ours 74.0 $\pm$ 3.7 2.7 $\pm$ 0.4 40.0 $\pm$ 3.2 6.3 $\pm$ 0.4 100 $\pm$ 0.0
IV-C Real Robots
To demonstrate the applicability of SkillWrapper for real-world robotic settings, we designed two sets of experiments with two robotic platforms: a Franka Emika Panda robot (Figures 3(a) and 3(b)) and a bimanual platform with two Kuka iiwa robots (Figures 3(c) and 3(d)). For both robot experiments, we assume that all skills are deterministic, which is a common assumption made by existing work [57, 22, 39, 4] (see Appendix -F for further details). We used an interaction budget of 15 steps per iteration for all methods requiring active data collection. The supplementary material contains video demonstrations of both experiments.
Generalization of SkillWrapper
In this setting, a tabletop Franka Emika Research 3 (Panda) manipulator has a skill set $\Omega$ consisting of five black-box skills: Pick, Place, Stack, Pour, and Wipe. The object set $\mathcal{O}$ contains five objects: a mug, a teapot, a plate, a bowl, and a sponge. The robot can pick up and place all objects except the mug and plate, pour ingredients from the teapot into the mug, and use the sponge to wipe the plate if it is dirty. To validate the generalization ability of the learned skill representations, we design three smaller training environments, where each environment contains a subset of $\mathcal{O}$, and thus only a subset of $\Omega$ is executable. Each training environment contains fewer than 10 possible states. After running SkillWrapper for exactly one iteration for each environment, we port the learned skill representations to the test environment that contains all objects in $\mathcal{O}$, which induces a state space of 34 possible abstract states. To quantify this generalization process, we similarly prepare an evaluation problem set that consists of five problems in the training environments, five problems in the test environment, and five Impossible problems across the two environments. The results are shown in Table II.
TABLE II: Results of Generalization Experiment
Method In-domain Generalization Impossible Solved % $\uparrow$ PB $\downarrow$ Solved % $\uparrow$ PB $\downarrow$ Expert Ops. 66.7 $\pm$ 9.4 3.3 $\pm$ 0.9 53.3 $\pm$ 9.4 5.3 $\pm$ 0.9 46.7 $\pm$ 0.0 ViLa [24] 46.7 $\pm$ 9.4 - 6.7 $\pm$ 9.4 - 6.7 $\pm$ 9.4 Random Exp. 0.0 $\pm$ 0.0 10.0 $\pm$ 0.0 0.0 $\pm$ 0.0 10.0 $\pm$ 0.0 60 $\pm$ 16.3 FM Invent 26.7 $\pm$ 18.9 7.3 $\pm$ 1.9 13.3 $\pm$ 9.4 8.7 $\pm$ 0.9 86.7 $\pm$ 9.4 No Heuristic 53.3 $\pm$ 9.4 4.7 $\pm$ 0.9 40.0 $\pm$ 16.3 6.0 $\pm$ 1.6 53.3 $\pm$ 9.4 Ours 76.7 $\pm$ 9.4 2.7 $\pm$ 0.9 60.0 $\pm$ 0.0 4.0 $\pm$ 0.0 66.7 $\pm$ 9.4
Learning Curve of SkillWrapper
In this bimanual manipulation setting, a robot with two Kuka iiwa arms is equipped with a skill set $\Omega$ containing six black-box skills: LeftArmPick, RightArmPick, Open, Scoop, Spread, and Drop. The object set $\mathcal{O}$ consists of three objects: a peanut butter jar, a knife, and a slice of bread. This robot can pick up the knife and jar, drop the knife, open the jar, scoop peanut butter with the knife, and spread it on the bread. Notably, this environment contains multiple dead ends, which would hinder the data gathering process. For example, the knife cannot be picked up again once dropped, the jar cannot be released once picked up, and the bread and knife cannot be cleaned once in contact with peanut butter. Moreover, the skills are highly interdependent. We designed the experiment in this way to investigate the learning process of SkillWrapper over several iterations. Again, we compare the performance against other baseline methods, ViLa and Expert Operators. An example of predictive truth value changes induced by a sequence is shown in Figure 5. We observe that performance improves as SkillWrapper progressively obtains more transition data and invents more predicates, finally surpassing the baselines. We plot the performance improvement over iterations in Figure 6.
(a) Initial
(b) InLeftGripper(Jar) $\rightarrow$ T
(c) Closed(Jar) $\rightarrow$ F
(d) InRightGripper(Knife) $\rightarrow$ T
(e) Coated(Knife) $\rightarrow$ T
(f) SpreadOn(Bread) $\rightarrow$ T
Figure 5: Illustration of a sequence of Bimanual Robot Skill Execution with Predicate Value Changes.
Discussion
Compared to the expert-constructed operators baseline, which serve as an approximate upper bound for performance without privileged simulator access, SkillWrapper demonstrates competitive accuracy while requiring only a small number of exploratory interactions. In particular, SkillWrapper exhibits strong generalization ability, indicating that the invented predicates and learned operators capture meaningful abstractions rather than overfitting to specific training traces. The gap between SkillWrapper and ViLa highlights the benefit of explicitly learning a symbolic model instead of relying solely on open-loop prompting, while the poor performance of Random Exploration underscores the importance of guided data collection in learning effective operators. Furthermore, the unstable performance of No Heuristic shows the importance of a sufficient predicate set and how SkillWrapper manages to reliably explore the state space, and the comparison against FM Invent shows how the inductive bias of our predicate invention algorithm facilitates task-level model learning. These findings suggest that SkillWrapper achieves a favorable trade-off between data efficiency and model soundness and completeness, which is crucial for scaling to larger domains. For brevity, we defer further detailed analysis of the baseline comparison to Appendix -H.
Our results demonstrate that SkillWrapper is effective in real robot settings: our method generalizes skill representations learned in restricted domains to richer environments and progressively improves in more challenging scenarios with irreversible actions and interdependent skills. By outperforming all baseline methods, SkillWrapper highlights the importance of predicate invention and iterative learning to scale symbolic representations to embodied tasks.
V Related Works
Our work uses pre-trained foundation models for learning symbolic representations of black-box robot skills useful to planning and close to human language and understanding. This work draws ideas from different fields of research such as model learning, abstraction learning, and task and motion planning (TAMP). Several methods have used foundation models (mainly LLMs) as high-level planners [1, 49, 13]. Several approaches have used foundation models as robot action models [6, 55] or to generate reward functions for robot tasks [66]. Concurrent work has also explored how representations can be learned directly from pixels [4]. Although these approaches show promising results for short-horizon single-skill problems, they fail to scale to complex long-horizon problems [31]. Lastly, multiple approaches [22, 39] have used foundation models to learn symbolic representations of robot skills, but require extensive feedback or prior knowledge from human experts. To the best of our knowledge, our work is the first to use a foundation model to automatically learn a human-interpretable symbolic representation of robot skills with theoretical guarantees and active data collection.
TAMP has long been used to solve complex robot tasks [11, 52, 17]. However, these approaches require symbolic models of robot skills for task planning. Various approaches have been developed to learn such symbolic models compatible with TAMP solvers from high-dimensional inputs [35, 57, 53]. Additionally, the abstract representations learned through these methods are not human-interpretable. In contrast, we explicitly design our approach to work with high-dimensional inputs and generate human-interpretable abstractions using pre-trained foundation models. We consider abstractions as human-interpretable if they are semantically meaningful and use informative language descriptions. SkillWrapper also connects to other domains in robotics, and the full related work can be found in Appendix -K.
Figure 6: Bimanual Kuka Scenario Results over five iterations with invented predicate and learned operator count of SkillWrapper. As the number of predicates and operators grows, SkillWrapper improves over the baseline methods.
VI Conclusion
We characterize important properties of a learned symbolic model and present the first known approach that employs off-the-shelf foundation models to invent symbolic representations for black-box skills of a robot while providing strong guarantees of soundness and completeness of the learned representations. By combining these theoretical guarantees with foundation model-driven data collection and predicate invention and evaluation, SkillWrapper facilitates principled predicate invention for task-level planning and produces interpretable operators directly usable by classical planners. Empirical results in simulation and on real robots demonstrate that SkillWrapper enables efficient long-horizon planning without hand-engineered abstractions, offering a principled path towards scalable skill representation learning.